Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.60 (given its target capital structure). Vandell has $8.10 million in debt that trades at par and pays an 7.8% interest rate. Vandell’s free cash flow (FCF0) is $2 million per year and is expected to grow at a constant rate of 4% a year. Both Vandell and Hastings pay a 40% combined federal and state tax rate. The risk-free rate of interest is 6% and the market risk premium is 6%. Hastings Corporation estimates that if it acquires Vandell Corporation, synergies will cause Vandell’s free cash flows to be $2.4 million, $3.0 million, $3.4 million, and $3.87 million at Years 1 through 4, respectively, after which the free cash flows will grow at a constant 4% rate. Hastings plans to assume Vandell’s $8.10 million in debt (which has an 7.8% interest rate) and raise additional debt financing at the time of the acquisition. Hastings estimates that interest payments will be $1.5 million each year for Years 1, 2, and 3. After Year 3, a target capital structure of 30% debt will be maintained. Interest at Year 4 will be $1.410 million, after which the interest and the tax shield will grow at 4%.

Indicate the range of possible prices that Hastings could bid for each share of Vandell common stock in an acquisition. Round your answers to the nearest cent. Do not round intermediate calculations
Answer #1

1. Calculation of Price of V Corporation :

Price = Free cash Flows (1+g)/ Ke-g

Calculation of Cost of Equity

Ke= Rf+Beta (Rm-Rf)

6+1.6 (6)

15.6%

Year Free Cash Flows PV DCF

1. 2 0.865 1.730

2. 2 0.748 1.497

3. 2 0.647 1.295

4. 2 0.560 1.120

Total 5.641

Price = 2(1.04)/ 15.6%-4%*2.821

= 50.58

Value of Equity Shares = Total Value - Debt

50.58+5.641- 8.1

48.12 Millions

Value per Share = 48.12/1

48.12/ Share

After Acquition of Vedels Corporation

Year Free Cash Flows PV DCF Interest addtional Net Cash flows

1. 2.4 0.865 2.08 1.5*0.5= 0.9 1.18

2. 3 0.748 2.24 0.9 1.34

3. 3.4 0.647 2.20 0.9 1.3

4. 3.87 0.560 2.17 1.140*0.6= 0.684 1.48

Total 8.69 5.31

Value of firm = 3.87 (1.04)/ 15.6% -4%* 2.821

(4.02 - 0.71136)/11.6% *2.821

80.58

Total Value = 80.58+5.31

85.88

Value of Equity = 85.88-8.1

77.78 Millions

Value per Share = 77.78/1

77.78

Therefore the Range is between 48 -78